OpenStudy (anonymous):

A rectangular plot of farmland will be bounded on one side by a river and on the other three sides by a single-strand electric fence. With 800m of wire at your disposal, what is the largest area you can enclose, and what are its dimensions.

OpenStudy (anonymous):

The largest area you can enclose is when the area is square. (800/3)^2=71111m^2

OpenStudy (anonymous):

its an optimization problem

OpenStudy (anonymous):

Hmm, you're right. It's only square if all four sides are bound by the fence. My mistake.

OpenStudy (anonymous):

isnt x+y=800 2x2y=P solve for y=800-x 2x(800-x)=P 1600x-2x^2=P f'(p)=1600-4x 4(400-x)=0 so x=400

OpenStudy (anonymous):

2x+y=800